# Asynchronous Message-Passing and Zeroth-Order Optimization Based Distributed Learning with a Use-Case in Resource Allocation in Communication Networks

### v2

@@ V Proposed Distributed learning
In accordance with Item 2, each agent $i$ can perform queries at time $t\in T_{q}^{i}$ by measuring its local reward function values and using the following expression:
@@
Here, $t\in\mathbb{N}$ represents the global time, $F_{i}$ is the reward function value (oracle) defined in Item 2, and the summations are over two minibatches of random variables $\bm{\xi}^{t,j}$ and $\bar{\bm{\xi}}^{{t,j}}$ for $j\in[B]$. While the oracle $F_{i}$ of agent $i$ relies on the (perturbed) parameters of its neighbors $j\in\mathcal{N}_{tx}^{i}$, agent $i$ does not require receiving the parameters $\bm{\theta}_{j}$. It only needs to query its own oracle, affected by $\bm{\theta}_{j}$ through their interaction.
It is essential to highlight that the query in (4) requires a certain level of synchrony among the agents. Specifically, the oracles $F_{i}(\bm{\theta}^{t},\bm{\xi}^{t,j})$ and $F_{i}(\bm{\theta}^{t}+\mu\bm{u}^{t},\bar{\bm{\xi}}^{t,j})$ necessitate the alignment of parameters $\bm{\theta}_{i}$ and $\bm{\theta}_{i}+\mu\bm{u}_{i}$ in time for all agents $i\in[m]$ to ensure the validity of the queries. This implies that the agents are required to perturb their local parameters synchronously. This requirement arises solely from employing zeroth-order oracles and can be lifted if higher-order oracles (e.g., function gradients) are utilized. Nonetheless, by II, agents still have the flexibility to asynchronously choose their preferred times $T_{u}^{i},T_{q}^{i}$ for updating, querying their oracles, and communicating with other agents, based on their heterogeneous communication and computational capabilities.
To update its parameter, each agent $i$ requires not only its own queries $\bar{r}_{i}^{t}$ but also queries from agents whose performance depends on $\bm{\theta}_{i}$, namely the agents in the set $\mathcal{N}^{i}_{rx}$ defined in (2). Parameter sharing is achieved through message-passing, which is susceptible to random (but bounded) communication delays. Therefore, each agent $i$ updates its parameter using potentially outdated queries received from its neighbors $\mathcal{N}^{i}_{rx}$. To establish the age of the queries, with a maximum $D^{\max}\geq 0$, consistent with the seminal work in [5], we introduce the timestamp $t-D^{\max}\leq\mathcal{T}_{i}^{j,t}\leq t$ defined as follows:
@@
It is evident that $\mathcal{T}_{i}^{i,t}=t,\forall i\in[m]$, $\mathcal{T}_{i}^{j,t}\in T_{q}^{j}$, and $0\leq t-\mathcal{T}_{i}^{j,t}\leq D^{\max}$ indicates how old agent $j$’s query stored in agent $i$’s memory is, with a maximum age of $D^{\max}$. An example of a message-passing architecture and the involved quantities are depicted in Fig. 1.
@@
In Algorithm 1, it is evident that utilizing zeroth-order oracles by each agent has three advantages. First, the proposed distributed learning approach can handle reward functions for which higher-order oracles (e.g., function gradients) may not be available or are expensive to calculate (e.g., in simulation-based optimizations or bandit optimizations [19]). Second, unlike federated learning or learning over graph settings where agents share local parameter or gradient vectors, employing zeroth-order oracles significantly reduces the communication bandwidth requirements in Algorithm 1 as agents only share scalars, by which the required gradients can be estimated (cf. 1.c and step 2 for message-passing; Items 2and 7 for gradient estimation). Third, sharing scalars ensures privacy-preserving cooperation and addresses the known problem of privacy leakage in federated learning settings [42, 43].
@@ VI-A Further definitions and assumptions
The functions defined in (11) represent the objective functions by which agent $i$ optimizes its local parameter $\bm{\theta}_{i}$. Based on (4), (9), Item 2, and Item 2, it can be observed (cf. further discussion in A.4 in Appendix) that the update $\bm{\tilde{s}}_{i}^{t}$ in Algorithm 1 satisfies
@@
In an ideal convergence scenario, we seek for $\Theta_{i}^{t}\to\bar{\Theta}^{t}_{i}$ (cf. VI.4), where the following holds for any $i\in[m]$:
@@
Ensuring a bounded delay is a crucial assumption in our convergence analysis (cf. VI.2). To establish such a guarantee, we impose the following assumption on the sets $T_{u}^{\ell}$ and $T_{q}^{\ell}$ in consistent with [5, Sec. 7]:
@@
Finally, we present the following required assumptions on the function gradients, oracles, and noise vectors:
@@
Assumption VI (noise vector).
@@
Items 1and 2 are standard in the stochastic optimization literature [44]. Additional Items 1and 2, however, is required specifically concerning the noise vector. The additional assumptions are necessary because the quantity $\bar{r}_{i}^{t}$ in (4) is obtained using different noise vectors for the perturbed and unperturbed functions. While most existing works in literature, e.g., [45, 44, 19, 24], assume that the oracle $F_{\ell}(\cdot,\bm{\xi})$ can be queried at two different points $\bm{\theta}$ and $\bm{\theta}^{\prime}$ in $\mathbb{R}^{n}$ with a single noise vector $\bm{\xi}$, this assumption does not hold in many practical applications, including the use-case discussed in Section VIII. In these applications, allowing different noise vectors $\bm{\xi}$ at different points $\bm{\theta}$ makes the implementation of zeroth-order optimization possible, as queries can be taken at separate time slots. We will further examine the impact of this practical limitation on the convergence of Algorithm 1 in the subsequent section (cf. VI.3), building upon studies of asynchronous parallel and distributed computations presented in [5].
@@ VI-B Main results
An important distinction in the complexity (17) lies in the presence of the last term in the variance term $\tilde{\sigma}^{2}$ which makes the tuning of variable $\mu$ challenging. In (17), it is evident that a smaller $\mu$ results in a smaller first term in the right-hand side; however, it also increases the variance $\tilde{\sigma}^{2}$. In VI.4, we will see that this results in convergence to a point closer to the stationary point but at a slower rate. It is essential to note that this term arises solely from the practical constraint that prohibits querying the oracle in two different points with the same noise vector. The following corollary establishes the iteration complexity when this constraint is lifted. The proof of this corollary is omitted here due to space limitations.
@@
Take s II, I, III, IVand V. In addition, it is assumed that the oracles $F_{\ell}(\cdot,\bm{\bm{\xi}}),\forall\ell\in[N]$ can be queried in any two points $\bm{\theta},\bm{\theta}^{\prime}\in\mathbb{R}^{n}$. Hence, the queries in (4) are performed by similar noise vectors in perturbed and unperturbed oracles for each $j$. Then for the iterates generated by Algorithm 1 the iteration complexity in (17) holds with the noise variance $\tilde{\sigma}^{2}\coloneqq 8(n+4)(\sigma^{2}+G^{2})+2{\mu^{2}}L^{2}(n+6)^{3}$.
@@
According to (18), a subsequence of $(\bm{\theta}^{t})_{t\in\mathbb{N}}$ converges to a neighborhood of some stationary point of $f$, and the size of this neighborhood is determined by the value of $\mu$. This convergence rate is consistent with existing works addressing zeroth-order nonconvex problems with synchronous updates, e.g., [45, 44, 19]. The rate $\mathcal{O}(\nicefrac{{1}}{{\sqrt{\bar{t}}}})$ can be established if the parameter $\mu$ also diminishes, i.e., is set according to $\mathcal{O}(\nicefrac{{1}}{{\sqrt{\bar{t}}}})$ [45, 44]. In addition, if the batch size increases proportionally to $\mathcal{O}(\sqrt{t})$, the rate $\mathcal{O}(\nicefrac{{1}}{{t}})$ is achieved [19]. If the assumption that the oracle $F_{\ell}(\cdot,\bm{\xi})$ can be queried at two different points $\bm{\theta}$ and $\bm{\theta}^{\prime}$ in $\mathbb{R}^{n}$ with a single noise vector $\bm{\xi}$ does not hold, which is the case in many practical applications, including the use-case considered in Section VIII, the convergence rate in (18) becomes $\mathcal{O}(\mu^{2})+\mathcal{O}(\nicefrac{{1}}{{(\mu^{2}\sqrt{\bar{t}})}})$, which can be concluded from the proof in Appendix B. Hence, in this case, a smaller $\mu$ guarantees convergence to a tighter neighborhood of some stationary point; however, the convergence rate becomes slower due to the last term.
@@ VII Implementation of the proposed distributed learning approach in RA problem
In this section, we explore the application of the proposed distributed learning in DL-based RA for communication networks, as an example (cf. Section II for a brief introduction). In the considered RA scenario, we aim to maximize the sum utility of agents (representing the global reward function) in the communication network, which can be formulated by (1). Since typical utility/loss functions in RA, such as data rate, communication latency, and appropriate variants that guarantee fairness, energy efficiency, etc., satisfy the assumptions considered in s I, Vand VI, the learning process can be addressed by the proposed learning approach in a distributed manner, benefiting from the advantages mentioned in Section I-B. It is important to note that for the considered example, in practice, each agent $i$, acting as a transmitter, can have a significant impact on the utility of a large number of agents due to cross-talk channels. For the purposes of this section, we assume that each agent $i$ influences the utility of all the other agents. Therefore, the neighbor sets defined in (2) are $\mathcal{N}^{i}_{rx}=[m]$ and $\mathcal{N}^{i}_{tx}=[m]$, necessitating a fully connected message-passing architecture where each agent communicates with all other agents, which is not always a practical solution. In this case, agent $i$ may choose to communicate with only a few agents (immediate neighbors) to gather the necessary information. Hence, the queries $\bar{r}_{\ell},\forall\ell\in[m]$ required by each agent $i$ are passed through intermediate agents and immediate neighbors (possibly with delays) to reach agent $i$. Consequently, in this application, it is necessary to redefine the message-passing architecture among agents and to establish new sets of neighbors.
@@ VII-A2 Communication between the agents
In the considered application, each agent affects the performance of all other agents through interaction (via cross-talk channels in the communication network). Therefore, it is important that each agent gathers the quantities $\bar{r}_{\ell}$ from all agents. Since each agent communicates only with its immediate neighbors defined in (19), it is crucial for each quantity to be associated with an agent ID as well. Hence, throughout the learning process, each agent $i$ stores and sends its query $\bar{r}_{i}^{t}$ along with a timestamp $t$ and the agent ID $i$. Similarly, the received quantities $\bar{r}_{\ell}^{\mathcal{T}_{i}^{\ell,t}}$ from the neighboring agents $j\in\bar{\mathcal{N}}^{i}_{rx}$ are also stored in the memory $\mathcal{B}_{i}$ along with their corresponding timestamps $\mathcal{T}_{i}^{\ell,t}$ and agent IDs $\ell\in[m]$. It is more practical if agent $i$ engages in communication with its neighbors $\bar{\mathcal{N}}^{i}_{rx}$ and $\bar{\mathcal{N}}^{i}_{tx}$ only at specific times of interest during the learning process. The time instants for communication are defined by the set $T^{i}_{tx}$ for agent $i$. At a given time $t\in T^{i}_{tx}$, agent $i$ transmits all the quantities $\bar{r}_{\ell}^{\mathcal{T}_{i}^{\ell,t}},\ell\in[m]$ (global information) stored in its memory to its neighbors $j\in\bar{\mathcal{N}}^{i}_{tx}$. The communication delay within the graph is random and bounded.
Since the graph $\mathcal{G}$ is connected, two agents may be connected through multiple paths. This means that agent $i$ may receive multiple quantities from agent $\ell\neq i$ with possibly different delays. In such cases, agent $i$ has the option to discard the older quantities and retain the most recent ones.
@@ VII-B Algorithm
The learning process requires the active participation of all agents to ensure that no agent is isolated and that each agent can transmit and receive the necessary information. This cooperative participation is guaranteed by VII. In addition, it is important to control the level of asynchrony for the newly introduced communication set $T^{\ell}_{tx}$ in order to bound the maximum delay $D^{\max}$ during the learning. This can be done, similar to III, by ensuring that each agent communicates at least every $D>0$ time instants.
@@
Since the underlying graph is not fully connected, each agent $i$ receives quantities $\bar{r}_{j},\forall j\in[m]$ with delays as they may pass through intermediate agents who may decide to transmit these quantities only at specific time instants of interest $T_{tx}^{\ell}$. The proposed asynchronous learning approach for DL-based RA is summarized in Algorithm 2.
@@
The sets $T_{q}^{i},T_{tx}^{i}$, and $T_{u}^{i}$ can be customized by agent $i$ based on factors such as its computational capacity, communication bandwidth, application layer, and other considerations. By specifying these sets, agent $i$ has the flexibility to determine when to query its reward, when to communicate with its neighbors, and when to update its local parameters.
It is important to note that the sets $T_{q}^{i},T_{tx}^{i}$, and $T_{u}^{i}$ are not required to be known to agent $i$ before the learning process begins. Rather, they are defined to facilitate the study of the proposed learning approach.

### v3

@@ V Proposed Distributed learning
In accordance with Item 2, each agent $i$ can perform queries at time $t\in T_{q}^{i}$ by measuring its local cost function values and using the following expression:
@@
Here, $t\in\mathbb{N}$ represents the global time, $F_{i}$ is the cost function value (oracle) defined in Item 2, and the summations are over two minibatches of sample vectors $\bm{\xi}^{t,j}$ and $\bar{\bm{\xi}}^{{t,j}}$ for $j\in[B]$. The queries in (4) with perturbed and unperturbed parameters can be performed in two different time slots, possibly using distinct sample vectors. While the oracle $F_{i}$ of agent $i$ relies on the (perturbed) parameters of its neighbors $j\in\mathcal{N}_{tx}^{i}$, agent $i$ does not require receiving the parameters $\bm{\theta}_{j}$. It only needs to query its own oracle $F_{i}$, affected by $\bm{\theta}_{j}$ through their interaction.
It is essential to highlight that the query in (4) requires a certain level of synchrony among the agents. Specifically, the oracles $F_{i}(\bm{\theta}^{t},\bm{\xi}^{t,j})$ and $F_{i}(\bm{\theta}^{t}+\mu\bm{u}^{t},\bar{\bm{\xi}}^{t,j})$ necessitate the alignment of parameters $\bm{\theta}_{i}$ and $\bm{\theta}_{i}+\mu\bm{u}_{i}$ in time for all agents $i\in[m]$ to ensure the validity of the queries. This implies that the agents should synchronously employ their perturbed and unperturbed local parameters, even if $t\notin T_{q}^{i}$. This requirement arises solely from employing zeroth-order oracles and can be lifted if higher-order oracles (e.g., function gradients) are utilized. Nonetheless, by II, agents still have the flexibility to asynchronously choose their preferred times $T_{u}^{i},T_{q}^{i}$ for updating, querying their oracles, and communicating with other agents, based on their heterogeneous communication and computational capabilities.
To update its parameter, each agent $i$ requires not only its own queries $\bar{r}_{i}^{t}$ but also queries from agents whose performance depends on $\bm{\theta}_{i}$, namely the agents in the set $\mathcal{N}^{i}_{rx}$ defined in (2). Parameter sharing is achieved through message-passing, which is susceptible to random (but bounded) communication delays. Therefore, each agent $i$ updates its parameter using potentially outdated queries received from its neighbors $\mathcal{N}^{i}_{rx}$. To establish the age of the queries, with a maximum $D^{\max}\geq 0$, consistent with the seminal work in [5], we introduce the timestamp $t-D^{\max}\leq\mathcal{T}_{i}^{\ell,t}\leq t$ defined as follows:
@@
It is evident that $\mathcal{T}_{i}^{i,t}=t,\forall i\in[m]$, $\mathcal{T}_{i}^{\ell,t}\in T_{q}^{\ell}$, and $0\leq t-\mathcal{T}_{i}^{\ell,t}\leq D^{\max}$ indicates how old agent $\ell$’s query stored in agent $i$’s memory is, with a maximum age of $D^{\max}$. An example of a message-passing architecture and the involved quantities are depicted in Fig. 1.
@@
In Algorithm 1, it is evident that utilizing zeroth-order oracles by each agent has three advantages. First, the proposed distributed learning approach can handle cost functions for which higher-order oracles (e.g., function gradients) may not be available or are expensive to calculate (e.g., in simulation-based optimizations or bandit optimizations [19]). Second, unlike federated learning or learning over graph settings where agents share local parameter or gradient vectors, employing zeroth-order oracles significantly reduces the communication bandwidth requirements in Algorithm 1 as agents only share scalars, by which the required gradients can be estimated (refer to 1.c and step 2 for message-passing; Items 2and 7 for gradient estimation). Third, sharing scalars adds a layer of privacy to distributed learning, mitigating the risk of privacy leakage, a known issue in federated learning settings. [42, 43].
@@ VI-A Further definitions and assumptions
The functions defined in (11) represent the objective functions by which agent $i$ optimizes its local parameter $\bm{\theta}_{i}$. Based on (4), (9), Item 2, and Item 2, it can be observed (refer to further discussion in A.4 in Appendix) that the update $\bm{\tilde{s}}_{i}^{t}$ in Algorithm 1 satisfies
@@
In an ideal convergence scenario, we seek for $\Theta_{i}^{t}\to\bar{\Theta}^{t}_{i}$ (refer to VI.4), where the following holds for any $i\in[m]$:
@@
Ensuring a bounded delay is a crucial assumption in our convergence analysis (refer to VI.2). To establish such a guarantee, we impose the following assumption on the sets $T_{u}^{\ell}$ and $T_{q}^{\ell}$ in consistent with [5, Sec. 7]:
@@
Finally, we present the following required assumptions on the function gradients, oracles, and sample vectors:
@@
Assumption VI (sample vector).
@@
Items 1and 2 are standard in the stochastic optimization literature [44]. Additional Items 1and 2, however, are required specifically concerning the sample vector. The additional assumptions are necessary because the quantity $\bar{r}_{i}^{t}$ in (4) is obtained using different sample vectors for the perturbed and unperturbed functions. While most existing works in literature, e.g., [45, 44, 19, 24], assume that the oracle $F_{\ell}(\cdot,\bm{\xi})$ can be queried at two different points $\bm{\theta}$ and $\bm{\theta}^{\prime}$ in $\mathbb{R}^{n}$ with a single sample vector $\bm{\xi}$, this assumption does not hold in many practical applications, including the use-case discussed in Section VIII. In these applications, allowing different sample vectors $\bm{\xi}$ at different points $\bm{\theta}$ makes the implementation of zeroth-order optimization possible, as queries can be taken at separate time slots. We will further examine the impact of this practical limitation on the convergence of Algorithm 1 in the subsequent section (refer to VI.3), building upon studies of asynchronous parallel and distributed computations presented in [5].
@@ VI-B Main results
An important distinction in the complexity (17) lies in the presence of the last term in the variance term $\tilde{\sigma}^{2}$ which makes the tuning of variable $\mu$ challenging. In (17), it is evident that a smaller $\mu$ results in a smaller first term in the right-hand side; however, it also increases the variance $\tilde{\sigma}^{2}$. In VI.4, we will see that this results in convergence to a point closer to the stationary point but at a slower rate. It is essential to note that this term arises solely from the practical constraint that prohibits querying the oracle in two different points with the same sample vector. The following corollary establishes the iteration complexity when this constraint is lifted. The proof of this corollary is omitted here due to space limitations.
@@
Take s II, I, III, IVand V. In addition, it is assumed that the oracles $F_{\ell}(\cdot,\bm{\xi}),\forall\ell\in[N]$ can be queried in any two points $\bm{\theta},\bm{\theta}^{\prime}\in\mathbb{R}^{n}$. Hence, the queries in (4) are performed by similar sample vectors in perturbed and unperturbed oracles for each sample $j$. Then for the iterates generated by Algorithm 1 the iteration complexity in (17) holds with the sample variance $\tilde{\sigma}^{2}\coloneqq 8(n+4)(\sigma^{2}+G^{2})+2{\mu^{2}}L^{2}(n+6)^{3}$.
@@
According to (18), a subsequence of $(\bm{\theta}^{t})_{t\in\mathbb{N}}$ converges to a neighborhood of some stationary point of $f$, and the size of this neighborhood is determined by the value of $\mu$. This convergence rate is consistent with existing works addressing zeroth-order nonconvex problems with synchronous updates, e.g., [45, 44, 19]. The rate $\mathcal{O}(\nicefrac{{1}}{{\sqrt{\bar{t}}}})$ can also be achieved if the parameter $\mu$ diminishes proportionally to $\mathcal{O}(\nicefrac{{1}}{{\sqrt{\bar{t}}}})$ [45, 44]. In addition, if the batch size increases proportionally to $\mathcal{O}(\sqrt{t})$, the rate $\mathcal{O}(\nicefrac{{1}}{{t}})$ is achieved [19]. If the assumption that the oracle $F_{\ell}(\cdot,\bm{\xi})$ can be queried at two different points $\bm{\theta}$ and $\bm{\theta}^{\prime}$ in $\mathbb{R}^{n}$ with a single sample vector $\bm{\xi}$ does not hold, which is the case in many practical applications, including the use-case considered in Section VIII, the convergence rate in (18) becomes $\mathcal{O}(\mu^{2})+\mathcal{O}(\nicefrac{{1}}{{(\mu^{2}\sqrt{\bar{t}})}})$, which can be concluded from the proof in Appendix B. Hence, in this case, a smaller $\mu$ guarantees convergence to a tighter neighborhood of some stationary point; however, the convergence rate becomes slower due to the last term.
@@ VII Implementation of the proposed distributed learning approach in RA problem
In this section, we explore the application of the proposed distributed learning in DL-based RA for communication networks, as an example (refer to Section II for a brief introduction). In the considered RA scenario, we aim to maximize the sum utility of agents (representing the global reward function) in the communication network, which can be formulated by (1). Since typical utility/loss functions in RA, such as data rate, communication latency, and appropriate variants that guarantee fairness, energy efficiency, etc., satisfy the assumptions considered in s I, Vand VI, the learning process can be addressed by the proposed learning approach in a distributed manner, benefiting from the advantages mentioned in Section I-B. It is important to note that for the considered example, in practice, each agent $i$, acting as a transmitter, can have a significant impact on the utility of a large number of agents due to cross-talk channels. For the purposes of this section, we assume that each agent $i$ influences the utility of all the other agents. Therefore, the neighbor sets defined in (2) are $\mathcal{N}^{i}_{rx}=[m]$ and $\mathcal{N}^{i}_{tx}=[m]$, necessitating a fully connected message-passing architecture where each agent communicates with all other agents, which is not always a practical solution. In this case, agent $i$ may choose to communicate with only a few agents (immediate neighbors) to gather the necessary information. Hence, the queries $\bar{r}_{\ell},\forall\ell\in[m]$ required by each agent $i$ are passed through intermediate agents and immediate neighbors (possibly with delays) to reach agent $i$. Consequently, in this application, it is necessary to redefine the message-passing architecture among agents and to establish new sets of neighbors.
@@ VII-A2 Communication between the agents
In the considered application, each agent affects the performance of all other agents through interaction (via cross-talk channels in the communication network). Therefore, it is important that each agent gathers the quantities $\bar{r}_{\ell}$ from all agents. It is more practical if agent $i$ engages in communication with its neighbors $\bar{\mathcal{N}}^{i}_{rx}$ and $\bar{\mathcal{N}}^{i}_{tx}$ only at specific times of interest during the learning process. The time instants for communication are defined by the set $T^{i}_{tx}$ for agent $i$. At a given time $t\in T^{i}_{tx}$, agent $i$ transmits all the quantities $\bar{r}_{\ell}^{\mathcal{T}_{i}^{\ell,t}},\ell\in[m]$ (global information) stored in its memory to its neighbors $j\in\bar{\mathcal{N}}^{i}_{tx}$. The communication delay within the graph is random and bounded.
Since the graph $\mathcal{G}$ is connected, two agents may be connected through multiple paths. This means that agent $i$ may receive multiple quantities from agent $\ell\neq i$ with possibly different delays. In such cases, agent $i$ has the option to discard the older quantities and retain the most recent ones. Hence, it is crucial for each quantity to be associated with an agent ID as well. Therefore, throughout the learning process, each agent $i$ stores and sends its query $\bar{r}_{i}^{t}$ along with a timestamp $t$ and the agent ID $i$. Similarly, the received quantities $\bar{r}_{\ell}^{\mathcal{T}_{i}^{\ell,t}}$ from the neighboring agents $j\in\bar{\mathcal{N}}^{i}_{rx}$ are also stored in the memory $\mathcal{B}_{i}$ along with their corresponding timestamps $\mathcal{T}_{i}^{\ell,t}$ and agent IDs $\ell\in[m]$.
@@ VII-B Algorithm
The learning process requires the active participation of all agents to ensure that no agent is isolated and that each agent can transmit and receive the necessary information. This cooperative participation is guaranteed by VII. In addition, it is important to control the level of asynchrony for the newly introduced communication instants set $T^{\ell}_{tx}$ in order to bound the maximum delay $D^{\max}$ during the learning. This can be done, similar to III, by ensuring that each agent communicates at least every $D>0$ time instants.
@@
Since the underlying graph is not fully connected, each agent $i$ receives quantities $\bar{r}_{\ell},\forall\ell\in[m]$ with delays as they may pass through intermediate agents who may decide to transmit these quantities only at specific time instants of interest $T_{tx}^{j}$. The proposed asynchronous learning approach for DL-based RA is summarized in Algorithm 2.
@@
The sets $T_{q}^{i},T_{tx}^{i}$, and $T_{u}^{i}$ can be customized by agent $i$ based on factors such as its computational capacity, communication bandwidth, application layer, and other considerations. By specifying these sets, agent $i$ has the flexibility to determine when to query its cost, when to communicate with its neighbors, and when to update its local parameters. It is important to note that the sets $T_{q}^{i},T_{tx}^{i}$, and $T_{u}^{i}$ are not required to be known to agent $i$ before the learning process begins. Rather, they are defined to facilitate the study of the proposed learning approach.